OpenStudy (anonymous):

identify which element is oxidized and which element is reduced in each of the following reaction and indicate the oxidation states of all atoms before and after the reaction. CH4(g) + 2O2(g) = CO2(g) + 2H2O(g). Zn(s) + 2HCL = ZnCl2(aq) + H2(g).

OpenStudy (anonymous):

Always remember OiL RiG. Oxidation is LOSS of electrons/hydrogen and gain of oxygen and Reduction is GAIN of electrons/hydrogen and loss of oxygen. For my classes, it's better to think in terms of electrons, but always remember about oxygen. That being said, in CH4 + 2O2 = CO2 + 2H2O The CH4 turns into CO2 by gaining oxygen and O2 turns into H2O by losing oxygen. The first is a reduction and the second is oxidation. For, Zn(s) + 2HCL = ZnCl2(aq) + H2(g). The HCl gains a hydrogen to become H2 (reduced), and the Zn gains Cl to become ZnCl2 (reduced). Think of the half reactions. Hope hat helps!